# RoToR: Towards More Reliable Responses for Order-Invariant Inputs

### v2

@@ Abstract
Mitigating positional bias of language models (LMs) for listwise inputs is a well-known and important problem (e.g., lost-in-the-middle). While zero-shot order-invariant LMs have been proposed to solve this issue, their success on practical listwise problems has been limited. In this work, as a first contribution, we identify and overcome two limitations to make zero-shot invariant LMs more practical: (1) training and inference distribution mismatch arising from modifying positional ID assignments to enforce invariance, and (2) failure to adapt to a mixture of order-invariant and sensitive inputs in practical listwise problems. Then, to overcome these issues we propose (1) RoToR, a zero-shot invariant LM for genuinely order-invariant inputs with minimal modifications of positional IDs, and (2) Selective Routing, an adaptive framework that handles both order-invariant and order-sensitive inputs in listwise tasks. On the Lost in the middle (LitM), Knowledge Graph QA (KGQA), and MMLU benchmarks, we show that RoToR with Selective Routing can effectively handle practical listwise input tasks in a zero-shot manner. [^1] https://github.com/soyoung97/RoToR
@@
Soyoung Yoon1thanks:   Work done during an internship at Channel Corporation.      Dongha Ahn1    Youngwon Lee1    Minkyu Jung2    HyungJoo Jang2    Seung-won Hwang1thanks:   Corresponding author. 1Seoul National University 2Channel Corporation {soyoung.yoon, seungwonh}@snu.ac.kr
@@ 1 Introduction
To overcome this, we propose a query-agnostic global sorting with circular arrangement for order-invariant positional ID assignment. Ours is named RoToR, inspired by the word rotary to express circular assignment, and also a palindrome, to reflect order invariance. Fig. 1c contrasts with PINE in Fig. 1c, where RoToR only needs a single global ordering (e.g., A->B->O->K->G) with no extra attention computation. The ordering of segments on suffix tokens remains in a fixed order, since it does not rely on their similarity to the query. Finally, we propose three different global sorting algorithms for RoToR, and demonstrate that they consistently outperform previous order-invariant models.
@@ 2.1 Positional bias of LLMs
Problem statement. Recent works on (zero-shot) retrieval augmented generation (RAG) with LLMs have found that the models exhibit unwanted bias on the ordering of the retrieved documents Chhabra et al. (2024). Widely known as the lost-in-the-middle problem Liu et al. (2024), many prior studies Chen et al. (2024); Gupta et al. (2024); Pezeshkpour and Hruschka (2023); Zhao et al. (2023); Zhou et al. (2024); Wei et al. (2024); Alzahrani et al. (2024); Zheng et al. (2024a) also investigate the importance of positional bias, extending the domain to structured knowledge grounding (SKG) tasks Zhao et al. (2023); Zhou et al. (2024) and multiple-choice questions Gupta et al. (2024) where changing the ordering of rows, schemas, or choices greatly degrades performance.
Considerations for decoder-only LMs. While successful approaches are presented to mitigate this issue for encoder-only Yang et al. (2022) and encoder-decoder Yen et al. (2024); Cai et al. (2023) models, they leave decoder-only models, which account for most of the current LLMs, for more consideration. In contrast to transformer encoders that use bidirectional attention which is invariant by nature Lee et al. (2019), transformer decoders use causal attention to learn causal relation signals, which is not invariant by nature Haviv et al. (2022a). Therefore, positional bias for decoder-only models is known to stem from both positional encoding and causal attention mask Yu et al. (2024); Wang et al. (2024) and is harder to mitigate.
@@ 2.2 Zero-shot order-invariance for LLMs
(Zero-shot) order-invariance. Recent line of works focused on achieving order-invariance by mechanistically altering both positional encoding and causal masking. While several works require training Junqing et al. (2023); Zhu et al. (2023), we focus on zero-shot approaches for practicality, namely PCW, Set-Based Prompting Ratner et al. (2023); McIlroy-Young et al. (2024), and PINE Wang et al. (2024), which we explain in detail at Sec. 3.1. Another line of works based on self-consistency try to mitigate positional bias simply by running inference multiple times with different orderings of contexts Zheng et al. (2024a). However, in principle, this requires evaluating $n!$ forward passes in total, enforcing Monte Carlo approximations (Tang et al., 2024). More recent work optimizes the number or passes (Lee et al., 2025b) with similar comprehensiveness (Hwang and Chang, 2007), or replaces with contrastive training objective Lee et al. (2025a). In contrast, our method guarantee invariance with a single forward pass, without requiring any approximations.
@@ 3 Methodology
Figure 2: For example input ‘‘Given Apple Banana Orange which one is most related to the color red? A: It is ...’’, demonstration of how PINE and RoToR (ours) modify the attention mask and positional IDs for each query-key combination, resulting in segment-wise order invariance. Assume each block represents a single token, and blocks are colored in positional ID assignment. White (no-block) area indicate masked attention.
@@ 3.1 Baseline: Order-invariant causal LMs
A more recent work, PINE (Wang et al., 2024) has addressed these issues through a bidirectional attention mechanism, by letting each segment attend to all other segments. However, to allow this within decoder-only models with causal attention and still achieve order invariance, PINE modifies the multi-head self-attention procedure to create an ‘illusion’: it treats each query segment as if it were the final segment in the input list, so that it can attend to all other segments as keys. For each query segment, the ordering among the key segments is determined by their attention scores (without positional embeddings, $\mathrm{Attn}_{\mathrm{NoPoS}}$), ensuring that segments with stronger relevance to the query appear closer. Meanwhile, the tokens within a query segment still follow causal ordering. Fig. 2 illustrates this with the input [T1‘‘Given’’, S1[‘‘Apple’’], S2[‘‘Ban’’, ‘‘ana’’], S3[‘‘Orange’’], T6‘‘which one’’, .. T10‘‘red?’’]. Prefix tokens ‘‘Given’’ and suffix tokens ‘‘which one .. red?’’ remain in their original positions and follow normal causal attention, equally attending to all segments. In contrast, the segments in the middle (S1 - S3) require the order-invariant mechanism. PINE dynamically re-assigns positional IDs depending on a token’s role as a query or a key, instead of using a single fixed positional ID. For Case 1: When “ana” (T4, S2) acts as a query, S2 is assigned the largest position IDs (5) so it can attend to all previous segments. Within S2, the tokens “ban” (4) and “ana” (5) still maintain their local causal order (i.e., smaller to larger positions). Case 2: When “ana” serves as a key, the placement of S2 depends on its $\mathrm{Attn}_{\mathrm{NoPoS}}$ scores with the query segment. If “banana” (S2) is more relevant to the query segment “apple” (S1) than “orange” (S3), (i.e., $\mathbf{Attn_{NoPos}}$(S1, S2) > $\mathbf{Attn_{NoPos}}$ (S1, S3)) S2 is placed nearer to S1, with “ban” and “ana” preserving their internal sequence. For prefix, suffix, and generated tokens, they do not participate in order invariance, so they are always placed at their standard causal positions. However, the same dynamic reordering applies to segment tokens (T2–T5) when computing attention scores of suffix (T6–T10) and the generated token (T11) as query.
@@ 3.2 RoToR: minimal OOD from positional ID assignments
While PINE achieves order-invariance by contextualization across segments, its query-specific ordering scheme introduces (1) significant train-test behavior discrepancy as well as (2) unnecessary complexity and numerical instability, which limits its scalability. During decoding with PINE, position IDs are assigned differently for every query token (each token in the suffix), decoder layer, and attention head, as the query-key attention score $\mathrm{Attn}_{\mathrm{NoPos}}$ determines the position IDs. This complexity introduces excessively frequent alterations on position IDs: As the base LM is trained with fixed positional IDs and causal masks, this causes hidden activations higher risk of out-of-distribution (OOD) for it to process properly. Moreover, ordering segments based on attention is computationally expensive and introduces numerical instability. As computing the attention value of one query segment requires computing the KV attention over every other number of segments, PINE invokes $\mathcal{O}(n^{2})$ cost overhead for each segment for input length $n$, which is further multiplied by the number of all combinations of layers, heads, and the number of suffix and generated tokens. Also, in practice, calculating attention without RoPE results in a very narrow range of values. bfloat16 numeric type lacks precision to distinguish these values, leading to non-determinism originating from several tied values. The outcome may depend on the initial ordering; to address these problems stemming from query-dependent ordering, we instead propose RoToR (Fig. 3), which uses one global ordering that is not a function of the initial ordering of segments (e.g., canonical ordering by lexical sorting) and assign IDs for tokens in different segments based on circular arrangement.
@@
Instead of re-computing the relative order of segments for each query, we reuse a globally shared single ordering, avoiding costly recomputation of numerically unstable attention scores. Moreover, this further reduces the gap between the LLM’s pretrained behavior and test-time behavior, as consistent position IDs are assigned across layers/heads/across suffix tokens. Global ordering allows to preserve the relative placement of segments, further closing the gap induced from introducing order invariance to causal LMs. For example, in Figure 3, due to the global ordering, segments $S_{5}$ and $S_{2}$ are always placed in adjacent positions with RoToR (right side), while it is not satisfied and constantly changed with PINE (left side). We consider three separate global sorting algorithm to be used in RoToR: (1) simple lexicographical sorting which can be obtained with minimal overhead based on tokenized sequence of segments, (2) using a pointwise reranker333castorini/monot5-base-msmarco-10k to score relevancy of each row with respect to the question, or (3) simple frequency-based sorting which normalizes token ids based on the inverse frequency of each token (Details at Fig. 6). Empirically, we find that using simple lexicographical sorting is sufficient for bringing improvements over PINE.
@@
By adopting global sorting, we can avoid extra attention scores computation and thus improve efficiency. While the cost with PINE to obtain hidden states is $\mathcal{O}(n^{2}d+nk\log{k})$ for input length $n$, hidden dimension $d$, and $k$ segments (Wang et al., 2024), our method with lexicographical sorting achieves $\mathcal{O}(nk\log{k})$ cost, being more efficient and faster (lightweight) than PINE. We empirically validate that our method is much faster than PINE as $k$ increases, at Tab. 4.
@@ 3.3 Selective Routing for handling order-sensitive inputs
Figure 4: Illustration of Selective Routing (Sec. A.5).
Since many practical benchmarks such as MMLU involves semi-invariant inputs, we propose a routing mechanism that uses the order-invariant model in conjunction with the standard causal model for further applicability. Our design is partly based on the finding from Wei et al. (2024) that there is correlation between task difficulty (which is in turn correlated with confidence values) and the model’s sensitivity to ordering. Selective Routing, illustrated in Fig. 4, combines confidence, the model output probability for the generated answer, from two different model versions—the original model and the order-invariant model—on the same input and choose a more confident answer. Both models first produce a maximum probability over possible answer tokens (e.g., A, B, C, D for MMLU) and a corresponding answer choice. We then compare the original model’s maximum probability, plus a bias term $\alpha$, to the invariant model’s maximum probability. If the original model’s adjusted score is higher, we take its answer; otherwise, the invariant model’s answer is chosen. $\alpha$ is a hyperparameter that controls how strongly the original model is favored, which was selected as 0.2 according to hyperparameter search on the validation subset (Appendix Sec. A.5).
@@ 4 Experiment setup
Llama-3.1-8B-Instruct Qwen1.5-4B-Chat Qwen1.5-7B-Chat
N = 30 N = 50 N = 30 N = 50 N = 30 N = 50
Method Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1
@@
Original 49.5 43.3 51.0 49.7 43.5 51.0 30.1 27.5 34.7 30.3 27.6 35.0 31.4 27.3 35.0 31.6 27.9 35.5
$\hookrightarrow$ stdev. ($\pm$) 0.07 0.14 0.17 0.34 0.28 0.46 0.41 0.34 0.43 0.26 0.24 0.35 0.26 0.28 0.29 0.40 0.56 0.42
PINE 51.8 45.2 52.8 51.8 45.3 52.7 31.5 28.7 35.6 31.5 28.7 35.3 32.3 28.8 35.7 31.7 28.2 35.7
$\hookrightarrow$ stdev. ($\pm$) 0.05 0.07 0.16 0.15 0.16 0.19 0.20 0.18 0.13 0.17 0.20 0.21 0.17 0.20 0.13 0.18 0.16 0.14
RoToR 52.8 46.2 53.8 52.7 45.9 53.5 31.8 28.8 35.5 32.5 29.6 36.1 34.2 29.9 37.7 34.2 30.1 38.0
$\hookrightarrow$ stdev. ($\pm$) 0.05 0.05 0.02 0.05 0.09 0.04 0.05 0.02 0.09 0.11 0.06 0.09 0.09 0.07 0.06 0.06 0.05 0.04
RoToR-MonoT5 51.6 45.0 52.6 52.2 45.2 52.8 32.4 29.2 36.3 32.3 29.4 35.9 33.0 28.8 36.5 32.8 28.8 36.5
$\hookrightarrow$ stdev. ($\pm$) 0.12 0.06 0.10 0.16 0.18 0.18 0.04 0.02 0.13 0.16 0.13 0.07 0.12 0.09 0.07 0.16 0.09 0.07
RoToR-Freq. 52.5 45.9 53.5 53.1 46.4 53.7 32.3 29.3 36.0 32.4 29.3 36.1 33.8 29.6 37.4 33.7 29.6 37.4
$\hookrightarrow$ stdev. ($\pm$) 0.10 0.15 0.11 0.02 0.07 0.03 0.13 0.16 0.09 0.09 0.04 0.06 0.04 0.00 0.09 0.04 0.16 0.22
Table 2: Results on the Mintaka (KGQA) dataset on different models, with standard deviation of the average scores with $\pm$. N refers to number of top-k segments per query.
@@
RoToR + S.R. (Optim.) 75.0 71.9 72.7 $\pm$ 1.0 61.8 60.1 61.1 $\pm$ 1.0 68.1 66.2 67.2 $\pm$ 0.7
Table 3: Improving applicability to general listwise tasks (MMLU, N=4) with Selective Routing (S.R), which includes both order-invariant and order-sensitive examples. RoToR with Selective Routing shows improved performance and stability across re-orderings of input, and its high upper-bound (Optim.) implies the potential for further improvements.

### v3

@@ Abstract
Mitigating positional bias of language models (LMs) for listwise inputs is a well-known and important problem (e.g., lost-in-the-middle). While zero-shot order-invariant LMs have been proposed to solve this issue, their success on practical listwise problems has been limited. In this work, as a first contribution, we identify and overcome two limitations to make zero-shot invariant LMs more practical: (1) training and inference distribution mismatch arising from modifying positional ID assignments to enforce invariance, and (2) failure to adapt to mixture of order-invariant and sensitive inputs in practical listwise problems. Then, to overcome these issues we propose (1) RoToR, a zero-shot invariant LM for genuinely order-invariant inputs with minimal modifications of positional IDs, and (2) Selective Routing, an adaptive framework that handles both order-invariant and order-sensitive inputs in listwise tasks. On the Lost in the middle (LitM), Knowledge Graph QA (KGQA), and MMLU benchmarks, we show that RoToR with Selective Routing can effectively handle practical listwise input tasks in a zero-shot manner. [^1] https://github.com/soyoung97/RoToR
@@
Soyoung Yoon1thanks:   Work done during an internship at Channel Corporation.      Dongha Ahn12    Youngwon Lee1    Minkyu Jung2    HyungJoo Jang2    Seung-won Hwang1thanks:   Corresponding author. 1Seoul National University 2Channel Corporation {soyoung.yoon, seungwonh}@snu.ac.kr
@@ 1 Introduction
To overcome this, we propose a query-agnostic global sorting with circular arrangement for order-invariant positional ID assignment. Ours is named RoToR, inspired by the word rotary to express circular assignment, and also a palindrome, to reflect order invariance.
Fig. 1c contrasts with PINE in Fig. 1c, where RoToR only needs a single global ordering (e.g., A->B->O->K->G) with no extra attention computation. The ordering of segments on suffix tokens remains in a fixed order, since it does not rely on their similarity to the query. Finally, we propose three different global sorting algorithms for RoToR, and demonstrate that they consistently outperform previous order-invariant models.
@@ 2.1 Positional bias of LLMs
Problem statement. Recent works on (zero-shot) retrieval augmented generation (RAG) with LLMs have found that the models exhibit unwanted bias on the ordering of the retrieved documents Chhabra et al. (2024). Widely known as the lost-in-the-middle problem Liu et al. (2024), many prior studies Chen et al. (2024); Gupta et al. (2024); Pezeshkpour and Hruschka (2023); Zhao et al. (2023); Zhou et al. (2024); Wei et al. (2024); Alzahrani et al. (2024); Zheng et al. (2024a) also investigate the impact of positional bias, extending the domain to structured knowledge grounding (SKG) tasks Zhao et al. (2023); Zhou et al. (2024) and multiple-choice questions Gupta et al. (2024) where changing the ordering of rows, schemas, or choices greatly degrades performance.
Considerations for decoder-only LMs. While successful approaches are presented to mitigate this issue for encoder-only Yang et al. (2022) and encoder-decoder Yen et al. (2024); Cai et al. (2023) models, they leave decoder-only models, which account for the current frontier LLMs, for more consideration. In contrast to transformer encoders that use bidirectional attention which is invariant by nature Lee et al. (2019), transformer decoders use causal attention to learn causal relation signals, which is not invariant by nature Haviv et al. (2022a). Therefore, positional bias for decoder-only models is known to stem from both positional encoding and causal attention mask Yu et al. (2024); Wang et al. (2024) and is harder to mitigate.
@@ 2.2 Zero-shot order-invariance for LLMs
(Zero-shot) order-invariance. Recent line of works focused on achieving order-invariance by mechanistically altering both positional encoding and causal masking. While several works require training Junqing et al. (2023); Zhu et al. (2023), we focus on zero-shot approaches for practicality, namely PCW Ratner et al. (2023), Set-Based Prompting McIlroy-Young et al. (2024), and PINE Wang et al. (2024), which we explain in detail at Sec. 3.1. Another line of works based on self-consistency try to mitigate positional bias simply by running inference multiple times with different orderings of contexts Zheng et al. (2024a). However, in principle, this requires evaluating $n!$ forward passes in total, enforcing Monte Carlo approximations (Tang et al., 2024). More recent work optimizes the number or passes (Lee et al., 2025b) with similar comprehensiveness (Hwang and Chang, 2007), or replaces with contrastive training objective Lee et al. (2025a). In contrast, our method guarantee invariance with a single forward pass, without requiring any approximations.
@@ 3 Methodology
Figure 2: Attention mask and positional ID modifications for segment-wise order invariance using example input ‘‘Given Apple Banana Orange...’’. Each block represents a single token arranged by positional ID assignment. In practice, tokens retain original positions but receive reassigned positional IDs as shown. White areas indicate masked attention. Key difference: PINE requires per-query sorting while RoToR reuses one global ordering across all queries.
@@ 3.1 Baseline: Order-invariant causal LMs
A more recent work, PINE (Wang et al., 2024) has addressed these issues through a bidirectional attention mechanism that allows each segment to attend to all other segments. To achieve this within decoder-only models while maintaining order invariance, PINE dynamically modifies positional IDs based on whether a token acts as a query or key in the attention computation.
The key insight is that PINE creates an “illusion” for each query segment: it assigns the query segment the largest positional IDs among all segments, enabling it to attend to all other segments bidirectionally. The ordering of key segments is then determined by their relevance scores computed without positional embeddings ($\mathrm{Attn}_{\mathrm{NoPoS}}$), ensuring that more relevant segments appear closer to the query.
Consider the example in Fig. 2 with input [T1‘Given’’, S1[Apple’’], S2[Ban’’, ana’’], S3[Orange’’], T6‘which one’’, .. T10‘red?’’]. The prefix token ‘Given’’ (T1) and suffix tokens ‘which one .. red?’’ (T6-T10) maintain their original positions and follow standard causal attention. For the segments S1-S3, PINE applies its order-invariant mechanism:
[…]
This dynamic reassignment occurs for every attention computation: each query token sees a different positional arrangement of the key tokens, determined by their relevance scores. Prefix, suffix, and generated tokens do not participate in this reordering and always maintain their standard causal positions. However, when these non-segment tokens act as queries, they still see the segments reordered by their relevance scores.
@@ 3.2 RoToR: minimal OOD from positional ID assignments
Challenges with PINE
While PINE achieves order-invariance by contextualization across segments, its query-specific ordering scheme introduces (1) significant train-test behavior discrepancy as well as (2) unnecessary complexity and numerical instability, which limits its scalability. During decoding with PINE, position IDs are assigned differently for every query token (each token in the suffix), decoder layer, and attention head, as the query-key attention score $\mathrm{Attn}_{\mathrm{NoPos}}$ determines the position IDs. This complexity introduces excessively frequent alterations on position IDs: As the base LM is trained with fixed positional IDs and causal masks, this causes hidden activations higher risk of out-of-distribution (OOD) for it to process properly. Moreover, ordering segments based on attention is computationally expensive and introduces numerical instability. As computing the attention value of one query segment requires computing the KV attention over every other number of segments, PINE invokes $\mathcal{O}(n^{2})$ cost overhead for each segment for input length $n$, which is further multiplied by the number of all combinations of layers, heads, and the number of suffix and generated tokens. Also, in practice, calculating attention without RoPE results in a very narrow range of values. bfloat16 numeric type lacks precision to distinguish these values, leading to non-determinism originating from several tied values. The outcome may then depend on the initial ordering.
Motivation & Theoretical Foundation
While investigating ways to overcome the limitations of PINE, our central goal is to preserve order invariance while minimizing the complexity of the re-assignment of positions. We reason that defining a single global ordering scheme, not necessarily relying on attention scores, and re-using them across all queries can solve the problems stemming from query-dependent ordering. A circular assignment of a global order seems as a practical solution. The idea of using global sorting to achieve ordering invariance has been studied in set/graph ML domain (Murphy et al., 2019a, b), but to the best of our knowledge, using circular assignment of the global ordering, and the application to pre-trained language models, are our novel contributions. As a result, we propose RoToR(Fig. 3), which uses one global ordering that is not a function of the initial ordering of segments (e.g., canonical ordering by lexical sorting) and assigns IDs for tokens in different segments based on circular arrangement.
@@
Instead of re-computing the relative order of segments for each query, we reuse a globally shared single ordering, avoiding costly recomputation of numerically unstable attention scores. Moreover, this further reduces the gap between the LLM’s pretrained behavior and test-time behavior, as consistent position IDs are assigned across layers/heads/across suffix tokens. Global ordering allows to preserve the relative placement of segments, further closing the gap induced from introducing order invariance to causal LMs. For example, in Figure 3, due to the global ordering, segments $S_{5}$ and $S_{2}$ are always placed in adjacent positions with RoToR (right side), while it is not satisfied and constantly changed with PINE (left side). We consider three separate global sorting algorithm to be used in RoToR: (1) simple lexicographical sorting which can be obtained with minimal overhead based on tokenized sequence of segments, (2) using a pointwise reranker Nogueira et al. (2020)333castorini/monot5-base-msmarco-10k to score relevancy of each row with respect to the question, or (3) simple frequency-based sorting which normalizes token ids based on the inverse frequency of each token (Details at Appendix Fig. 6). Empirically, we find that using simple lexicographical sorting is sufficient to obtain improvements over PINE.
@@
We report only operations executed beyond vanilla self-attention cost $O(n^{2}d)$, where $n$ is total input length, $d$ is hidden dimension, and $k$ is the number of segments. PINE requires two additional operations: (1) computing attention scores without rotary position embeddings ($\mathcal{O}(n^{2}d)$) and (2) sorting $k$ segments for each query token ($\mathcal{O}(nk\log k)$), totaling $\bm{\mathcal{O}(n^{2}d+nk\log k)}$ (Wang et al., 2024) [^4] The PINE paper reports $\mathcal{O}(nk\log k)$ by absorbing the $\mathcal{O}(n^{2}d)$ term into baseline; we expose it explicitly for fair comparison.. In contrast, our lexicographical sorting requires only a single global sort of $k$ segments ($\mathcal{O}(k\log k)$), each with length $\mathcal{O}(n)$, achieving $\bm{\mathcal{O}(nk\log k)}$ and eliminating the expensive $\mathcal{O}(n^{2}d)$ term entirely. This can be further optimized to $\bm{\mathcal{O}(nk)}$ using radix sort. [^5] https://en.wikipedia.org/wiki/Radix_sort We empirically validate significantly faster performance than PINE as $k$ increases (Tab. 4).
@@ 3.3 Selective Routing for handling order-sensitive inputs
Figure 4: Illustration of Selective Routing (Sec. E).
Since many practical benchmarks such as MMLU involves semi-invariant inputs, we propose a routing mechanism that uses the order-invariant model in conjunction with the standard causal model for further applicability. Our design is partly based on the finding from Wei et al. (2024) that there is correlation between task difficulty (which is in turn correlated with confidence values) and the model’s sensitivity to ordering. Selective Routing, illustrated in Fig. 4, combines confidence, the model output probability for the generated answer, from two different model versions—the original model and the order-invariant model—on the same input and choose a more confident answer. Both models first produce a maximum probability over possible answer tokens (e.g., A, B, C, D for MMLU) and a corresponding answer choice. We then compare the original model’s maximum probability, plus a bias term $\alpha$, to the invariant model’s maximum probability. If the original model’s adjusted score is higher, we take its answer; otherwise, the invariant model’s answer is chosen. $\alpha$ is a hyperparameter that controls how strongly the original model is favored, which was selected as 0.2 according to hyperparameter search on the validation subset (Appendix Sec. E).
@@ 4 Experiment setup
Llama 3.1-70B-Instruct
Original 66.2 65.7 65.7 65.2 64.3 65.0 66.2 64.8
PINE 67.9 67.8 67.5 65.9 65.7 65.9 65.8 65.5
RoToR 69.6 69.5 69.3 67.6 67.8 67.8 67.7 67.9
RoToR-MonoT5 68.9 69.0 68.8 67.5 67.5 67.7 67.5 67.6
@@
Llama-3.1-8B-Instr. Llama-3.1-70B-Instr. Qwen1.5-4B-Chat Qwen1.5-7B-Chat Qwen1.5-72B-Chat
Method Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1
$N\!=\!30$
@@
Original 50.2 44.0 51.9 61.6 57.7 63.6 30.7 27.9 34.9 31.5 27.8 35.4 41.4 37.7 43.7
PINE 51.5 45.0 52.6 63.1 58.7 64.8 31.6 28.7 35.6 32.3 28.8 36.4 46.7 42.9 49.0
RoToR 53.1 46.5 54.1 63.6 59.1 65.2 32.0 29.0 35.7 34.3 29.8 37.7 47.5 43.2 49.2
RoToR-MonoT5 51.6 45.0 52.5 32.3 29.1 36.2 32.9 28.4 36.3
RoToR-Freq. 52.6 46.1 53.7 32.3 29.2 36.0 33.7 29.5 37.2
@@
Original 49.5 43.3 51.0 62.1 57.8 64.0 30.1 27.5 34.7 31.4 27.3 35.0 41.0 37.6 43.6
$\hookrightarrow$ stdev. ($\pm$) 0.07 / 0.14 / 0.17 0.37 / 0.40 / 0.27 0.41 / 0.34 / 0.43 0.26 / 0.28 / 0.29 0.75 / 0.40 / 0.33
PINE 51.8 45.2 52.8 63.3 58.8 64.9 31.5 28.7 35.6 32.3 28.8 35.7 46.9 43.3 49.2
$\hookrightarrow$ stdev. ($\pm$) 0.05 / 0.07 / 0.16 0.13 / 0.04 / 0.10 0.20 / 0.18 / 0.13 0.17 / 0.20 / 0.13 0.18 / 0.20 / 0.20
RoToR 52.8 46.2 53.8 63.5 59.1 65.3 31.8 28.8 35.5 34.2 29.9 37.7 47.4 43.1 49.1
$\hookrightarrow$ stdev. ($\pm$) 0.05 / 0.05 / 0.02 0.11 / 0.07 / 0.08 0.05 / 0.02 / 0.09 0.09 / 0.07 / 0.06 0.06 / 0.04 / 0.07
RoToR-MonoT5 51.6 45.0 52.6 32.4 29.2 36.3 33.0 28.8 36.5
$\hookrightarrow$ stdev. ($\pm$) 0.12 / 0.06 / 0.10 0.04 / 0.02 / 0.13 0.12 / 0.09 / 0.07
RoToR-Freq. 52.5 45.9 53.5 32.3 29.3 36.0 33.8 29.6 37.4
$\hookrightarrow$ stdev. ($\pm$) 0.10 / 0.15 / 0.11 0.13 / 0.16 / 0.09 0.04 / 0.00 / 0.09
$N\!=\!50$
Initial, no shuffling of segments
Original 50.0 44.0 51.7 62.6 58.5 64.5 31.6 28.6 35.8 31.7 28.0 35.7 42.1 38.7 44.5
PINE 51.6 45.1 52.6 64.1 59.8 65.8 31.6 28.8 35.3 32.0 28.5 35.9 48.0 44.1 49.9
RoToR 52.9 46.0 53.6 64.6 60.0 66.2 32.7 29.6 36.2 34.3 30.1 38.0 48.4 44.3 50.3
RoToR-MonoT5 52.4 45.4 52.8 32.3 29.3 35.9 32.9 28.9 36.6
RoToR-Freq. 53.1 46.4 53.7 32.3 29.2 36.1 33.5 29.5 37.2
After shuffling segments, averaged over 3 seeds
Original 49.7 43.5 51.0 62.8 58.5 64.5 30.3 27.6 35.0 31.6 27.9 35.5 42.1 38.9 44.7
$\hookrightarrow$ stdev. ($\pm$) 0.34 / 0.28 / 0.46 0.29 / 0.28 / 0.05 0.26 / 0.24 / 0.35 0.40 / 0.56 / 0.42 0.30 / 0.40 / 0.35
PINE 51.8 45.3 52.7 64.3 59.8 65.9 31.5 28.7 35.3 31.7 28.2 35.7 48.0 44.3 50.0
$\hookrightarrow$ stdev. ($\pm$) 0.15 / 0.16 / 0.19 0.16 / 0.15 / 0.14 0.17 / 0.20 / 0.21 0.18 / 0.16 / 0.14 0.02 / 0.04 / 0.05
RoToR 52.7 45.9 53.5 64.5 60.0 66.1 32.5 29.6 36.1 34.2 30.1 38.0 48.3 44.3 50.3
$\hookrightarrow$ stdev. ($\pm$) 0.05 / 0.09 / 0.04 0.02 / 0.02 / 0.01 0.11 / 0.06 / 0.09 0.06 / 0.05 / 0.04 0.05 / 0.09 / 0.05
RoToR-MonoT5 52.2 45.2 52.8 32.3 29.4 35.9 32.8 28.8 36.5
$\hookrightarrow$ stdev. ($\pm$) 0.16 / 0.18 / 0.18 0.16 / 0.13 / 0.07 0.16 / 0.09 / 0.07
RoToR-Freq. 53.1 46.4 53.7 32.4 29.3 36.1 33.7 29.6 37.4
$\hookrightarrow$ stdev. ($\pm$) 0.02 / 0.07 / 0.03 0.09 / 0.04 / 0.06 0.04 / 0.16 / 0.22
Table 2: Mintaka (KGQA) results, with Initial and After-shuffle settings, across different model parameter size and backbones. $N$ refers to the number of top-k segments per query. Rows with $\hookrightarrow$ stdev.” report standard deviation over 3 seeds.
@@
RoToR + S.R. (Oracle) 75.0 71.9 72.7 $\pm$ 1.0 61.8 60.1 61.1 $\pm$ 1.0 68.1 66.2 67.2 $\pm$ 0.7
Table 3: Improving applicability to general listwise tasks (MMLU, N=4) with Selective Routing (S.R), which includes both order-invariant and order-sensitive examples. Init./Rev. refer to original/reversed orderings, Avg. is the average selection ratio across all (4!-1) re-orderings with standard deviation. S.R (Oracle) represents the upper bound with perfect routing accuracy. RoToR with Selective Routing shows improved performance and stability across input re-orderings.